# CoLoRA: Continuous low-rank adaptation for reduced implicit neural modeling of parameterized partial differential equations

### v1

@@ Appendix B Details on numerical experiments
[…]
All numerical experiments were implemented in Python with JAX [14] with just-in-time compilation enabled. All benchmarks were run on a single NVIDIA RTX-8000 GPU.
@@ Appendix C Description of full order models (FOMs)
In order to ensure a fair comparison in terms of runtime between CoLoRA and the FOMs, we implement all FOMs in JAX [14] with just-in-time compilation.
@@ C.2 Burgers’
For the Burgers’ full order model we follow the full order model described [94]. This uses finite differences to compute the spatial derivatives and uses a fixed-time step implicit method with Newton iterations for time integration. For the full order model benchmark we choose a $1024\times 1024$ spatial grid.
@@ C.3 Rotating Detonating Engine
The implementation for the RDE full order model follows [89] which uses finite differences to compute the spatial derivatives and use a fixed-time step implicit method with Newton iterations for time integration.
@@ Appendix D Pre-training and architecture details
The width of all layers is 25 except the last whose width must be 1 in order to output a scalar field. In the case of the RDE example and the 2D Wave example given in [96] the last layer is of width 2 in order to output a field for each variable in the equation. The hyper-network $h$ is a multilayer perceptron of depth 3 which also uses swish nonlinear activation function. The width of each layer is 15, except the last layer whose width is $q$, the dimension of the online parameter vector $\boldsymbol{\phi}(t,\boldsymbol{\mu})$.
@@ D.1 Periodic $P$ layer
where $a,c,b\in\mathbb{R}^{d}$ are additionally part of the offline parameters $\boldsymbol{\theta}$. The only exception is in 1D Inviscid Burgers’ given in [19] which does not have periodic boundary conditions. Here we simply replace $P$ with another $\mathcal{C}$ layer. In this case the boundary are loosely enforced via pretraining.
@@ D.3 Pre-training
In pre-training for all our benchmark problems (Vlasov, Burgers’, and RDE) we minimize (6) using an Adam optimizer [51] with the following hyper-parameters,
@@
For the results given in Table 1 for the 1D Inviscid Burgers’ and 2D Wave problems, we use $250,000$ iterations, with all other hyper-parameters kept the same.
@@ Appendix E F-FNO experiments
For implementation of the F-FNO we use the code base given in the original paper [92] while keeping the modification that we make minimal. We use their largest architecture which is 24 layers deep as this was shown to give the best possible performance on their benchmarks. This was obtained via a grid sweep of the number of layers and time step size for the F-FNO. Additionally we give our $\boldsymbol{\mu}$ as input to their network. We train over 100 epochs as in their implementation. In order to give the F-FNO the best possible performance, the error reported is from the best possible checkpoint over all the epochs. All other hyper-parameters we set according to their implementation.
@@ Appendix F Linear projection as comparison with respect to best approximation error
In order to compute the optimal linear projection we assemble the training and test data into two snapshot matrices. We then compute the singular value decomposition of the training snapshot matrix and build a projection matrix from the top $n$ left singular values where $n$ is the reduced dimension. We then use this projection matrix to project the test data into the reduce space and then project back up into the full space using the transpose of the projection matrix. We then measure the relative error between the resulting project test data and the original test data. This value gives the optimal linear approximation error. For additional details see [54].
@@ Appendix H Neural Galerkin computational procedure
which is then discretized in time and solved for the corresponding trajectory of latent states $\boldsymbol{\phi}(t_{1},\boldsymbol{\mu}),\dots,\boldsymbol{\phi}(t_{K},\boldsymbol{\mu})\in\mathbb{R}^{q}$ at the time steps $t_{1}<\dots<t_{K}$. We refer to [15, 10] for details on this computational approach.

### v2

@@ Appendix B Details on numerical experiments
Refer to captionCoLoRA-EQRefer to captionCoLoRA-DRefer to captionF-FNO0.000.00{0.00}0.000.050.05{0.05}0.050.100.10{0.10}0.10Refer to captionBurgers’
Refer to captionCoLoRA-EQRefer to captionCoLoRA-DRefer to captionF-FNO0.000.00{0.00}0.000.050.05{0.05}0.050.100.10{0.10}0.10Refer to captionVlasov
@@
All numerical experiments were implemented in Python with JAX (Bradbury et al., 2018) with just-in-time compilation enabled. All benchmarks were run on a single NVIDIA RTX-8000 GPU.
@@ Appendix C Description of full order models (FOMs)
In order to ensure a fair comparison in terms of runtime between CoLoRA and the FOMs, we implement all FOMs in JAX (Bradbury et al., 2018) with just-in-time compilation.
@@ C.2 Burgers’
For the Burgers’ full order model we follow the full order model described (Wang et al., 2016). This uses finite differences to compute the spatial derivatives and uses a fixed-time step implicit method with Newton iterations for time integration. For the full order model benchmark we choose a $1024\times 1024$ spatial grid.
@@ C.3 Rotating Detonating Engine
The implementation for the RDE full order model follows (Singh et al., 2023) which uses finite differences to compute the spatial derivatives and use a fixed-time step implicit method with Newton iterations for time integration.
@@ Appendix D Pre-training and architecture details
The width of all layers is 25 except the last whose width must be 1 in order to output a scalar field. The only larger network is used in the 3D spherical shallow water example where the width is 128. In the case of the RDE example and the 2D Wave example given in Yin et al. (2023) the last layer is of width 2 in order to output a field for each variable in the equation. The hyper-network $h$ is a multilayer perceptron of depth 3 which also uses swish nonlinear activation function. The width of each layer is 15, except the last layer whose width is $q$, the dimension of the online parameter vector $\boldsymbol{\phi}(t,\boldsymbol{\mu})$.
@@ D.1 Periodic $P$ layer
where $a,c,b\in\mathbb{R}^{d}$ are additionally part of the offline parameters $\boldsymbol{\theta}$. The only exception is in 1D Inviscid Burgers’ given in (Chen et al., 2023b) which does not have periodic boundary conditions. Here we simply replace $P$ with another $\mathcal{C}$ layer. In this case the boundary are loosely enforced via pretraining.
@@ D.3 Pre-training
In pre-training for all our benchmark problems (Vlasov, Burgers’, and RDE) we minimize (6) using an Adam optimizer (Kingma & Ba, 2015) with the following hyper-parameters,
@@
For the results given in Table 1 for the 2D Wave and 3D Shallow Water problems, we use $250,000$ and $2,000,000$ iterations respectively, with all other hyper-parameters kept the same.
@@ Appendix E F-FNO experiments
For implementation of the F-FNO we use the code base given in the original paper (Tran et al., 2023) while keeping the modification that we make minimal. We use their largest architecture which is 24 layers deep as this was shown to give the best possible performance on their benchmarks. This was obtained via a grid sweep of the number of layers and time step size for the F-FNO. Additionally we give our $\boldsymbol{\mu}$ as input to their network. We train over 100 epochs as in their implementation. In order to give the F-FNO the best possible performance, the error reported is from the best possible checkpoint over all the epochs. All other hyper-parameters we set according to their implementation.
We provide additional results of the experiments in Table 2.
Method Vlasov (rel. err.) 2D Burgers (rel. err.) RDE (rel. err.)
High Data (100 Trajectories)
F-FNO 8.57e-3 5.11e-3 2.21e-3
CoLoRA-EQ 1.58e-3 2.27e-3 1.49e-3
CoLoRA-D 9.87e-4 4.96e-4 2.05e-4
Low Data (10 Trajectories)
F-FNO 7.48e-2 2.40e-2 5.69e-3
CoLoRA-EQ 2.73e-3 3.99e-3 1.79e-3
CoLoRA-D 2.37e-3 1.76e-3 4.47e-4
Table 2: Detailed results of F-FNO data efficiency experiment
@@ Appendix F Linear projection as comparison with respect to best approximation error
In order to compute the optimal linear projection we assemble the training and test data into two snapshot matrices. We then compute the singular value decomposition of the training snapshot matrix and build a projection matrix from the top $n$ left singular values where $n$ is the reduced dimension. We then use this projection matrix to project the test data into the reduce space and then project back up into the full space using the transpose of the projection matrix. We then measure the relative error between the resulting project test data and the original test data. This value gives the optimal linear approximation error. For additional details see (Kramer et al., 2024).
@@ Appendix H Neural Galerkin computational procedure
which is then discretized in time and solved for the corresponding trajectory of latent states $\boldsymbol{\phi}(t_{1},\boldsymbol{\mu}),\dots,\boldsymbol{\phi}(t_{K},\boldsymbol{\mu})\in\mathbb{R}^{q}$ at the time steps $t_{1}<\dots<t_{K}$. We refer to (Bruna et al., 2024; Berman & Peherstorfer, 2023) for details on this computational approach.